The perimeter of a rectangle is 74 feet. The width of the rectangle is three feet less than the length. What is the length and the width?
Answer by AAfter.Com(18)

Given, 2(l + b) = 74, l and b being length and breadth.
=> l + b = 37
Also b = l - 3
Hence, l + l - 3 = 37
=> 2l = 40
=> l = 20
b = l - 3 = 20 - 3 = 17
Hence, length and breadth are 20 and 17 units.